OpenStudy (anonymous):

8a - 24 + 3a = 6a + 1

OpenStudy (photon336):

@moonlight123 the first step yo need to do is to combine like terms do you know what that is?

OpenStudy (kendricklamar2014):

1. Simplify: 8a−24+3a=6a+1 becomes 8a+−24+3a=6a+1

OpenStudy (kendricklamar2014):

2. Combine Like Terms to get: (8a+3a)+(−24)=6a+1

OpenStudy (kendricklamar2014):

Which equals: 11a−24=6a+1

OpenStudy (kendricklamar2014):

3. Subtract 6a from both sides: 11a−24−6a=6a+1−6a which equals: 5a−24=1

OpenStudy (kendricklamar2014):

4. Add 24 to both sides: 5a−24+24=1+24 which equals: 5a=25

OpenStudy (anonymous):

ok =ay thank you

OpenStudy (kendricklamar2014):

5. Divide both sides by 5: 5a/5 = 25/5 which equals: a=5

OpenStudy (kendricklamar2014):

Your Welcome